OpenStudy (anonymous):

the third term of an arithmetic is 21, and the eigth term is 56. the first term is...

OpenStudy (anonymous):

2 equations 21=a+2d 56=a+7d

OpenStudy (anonymous):

use this for more reference . i am busy sorry contact me for help though ....XD http://openstudy.com/study#/updates/504cef62e4b02980ad8101c4

OpenStudy (anonymous):

the first term is 7

OpenStudy (anonymous):

@trina1478

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

lets get a notation and everything I will denote (that all mathematicians denote) straight. notation read as meaning \(\large\color{black}{ \displaystyle a_1 }\) a of 1 the first term in a sequence \(\large\color{black}{ \displaystyle a_2 }\) a of 2 the second term in a sequence \(\large\color{black}{ \displaystyle a_3 }\) a of 3 the third term in a sequence and so forth... and IN GENERAL. \(\large\color{black}{ \displaystyle a_n }\) a of n any \(\rm n\) th term in a sequence

OpenStudy (solomonzelman):

Now, in arithmetic sequence, all terms follow a certain pattern. You always add or subtract the same number to get to the next term. We refer to this number that you add or subtract (in an arithmetic sequence, to get to the next term) using a letter "d" (because it id the \(\rm d\)ifference between the terms)

OpenStudy (anonymous):

gice me medal i answered correctly